Light bulbs filled with an inert gas, with the electric off at T = 20 C the pressure is 115 KPa. The light bulb is turned on and the temperature rises to 70 C, determine the new pressure within the bulb.

a. 135 kPa b. 171 kPa c. 129 kPa d. 402 kPa
Answer #1

  P/T = constant for the fixed mass of gas inside the bulb

P1/T1 = P2/T2 .. [1=bulb off, 2=bulb on]
P2 = (T2/T1)*P1

TI = (273+20) = 293K
T2 = (273+70) = 343K

P2 = (343/293)*115kPa .. .. ►P2 = 134.62 kPa or 135 kpa